For a particular event, 362 tickets were sold for a total of $1912. if students paid $4 per ticket and non-students paid $6 per ticket, how many student tickets are sold

Answer by TimothyLamb(4379)

x = number of student tickets
y = number of non-student tickets
---
x + y = 362
4x + 6y = 1912
---
put the system of linear equations into standard form
---
x + y = 362
4x + 6y = 1912
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = number of student tickets = 130
y = number of non-student tickets = 232
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---